solve. x-2/x+1 greater or equal to 3
thank you

Answer by jai_kos(139)

Solution:
x-2 /x+1 > = 3
Multiply both sides by (x+1), we have
x -2 > = 3 (x+ 1)
x - 2 >= 3x + 3
Group all the x terms to one side, we get
x -3x > = 3 + 2
-2x > =5
Divide the above equation by 2 we get,
-x>= 5/2
When we change the sign in inequalites ,we also change the symbols.
x < = 5/2